# Large diversity of magnetic phases in two-dimensional magnets with spin-orbit coupling and superconductivity

### v1

@@ I Introduction
There is a vast versatility of magnetic phases in the physics of interfaces. Among ferromagnetic and antiferromagnetic spin orientations in the ground state, recently noncollinear phases including spirals, multi-q states [1, 2, 3], canted AFMs [4], bubbles, and topological magnetism including skyrmions [5] and merons [6], as well as further exotic phases like altermagnets [7] have been theoretically and experimentally discovered. Magnetic phases form a cornerstone of today’s and future storage technology [8, 9, 10] and are fundamentally interesting because of exotic classical and quantum quasiparticle excitations that could be hosted in frustrated or correlated magnets [11, 12]. For modeling purposes and for detecting yet novel magnetic phases, the question arises which conceptionally simple models host a large number of magnetic phases that can be tuned by experimentally accessible parameters. Particularly interesting in this regard are systems combining magnetism and superconductivity. While numerous studies of one-dimensional magnetic systems in proximity to s-wave superconductors exist, motivated by the interest in Majorana physics [13, 14, 15, 16], the data on phase diagrams of their two-dimensional counterparts have been very limited up to now [17, 18]. Ground state studies on quantum and classical magnetic moments coupled to itinerant electrons found evidence of chiral order [19], topological electronic order [20, 21], and multi-spin interactions [22] in frustrated triangular lattices without superconductivity. Motivated by recent studies on superconducting 1D systems [13], we therefore consider a two-dimensional square lattice tight-binding model with local magnetic fields, Rashba spin-orbit coupling, and superconductivity. Our model is effectively describing localized classical magnetic moments that are coupled by the itinerant electrons of a surface state, not specifically bound to any parameter regime. The knowledge of possible ground states in those complex structures is required for several reasons. First, exotic topological states have been found in 2D itinerant magnets in proximity to a superconductor (SC) [17, 18]. Second, interesting multi-spin interactions can be present in such systems [23, 1] that might lead to unexpected magnetic and electronic phases. Due to big datasets in the study of magnetic systems, machine learning has been applied extensively in this field, which we adopt as well. Corresponding techniques have especially been employed to predict the parameters of the magnetic Hamiltonian [24], to retrieve the topological charge from dynamic chiral magnetic structures [25], and to construct phase diagrams for skyrmionic systems [26, 27]. Additionally, unsupervised learning has found a variety of applications in the field of microscopic magnetism [28, 29, 30] and the identification of topological phases [31, 32].
@@
In this manuscript, we analyze the ground states of magnetic moments coupled by a spinful tight-binding model with proximity induced s-wave superconductivity, local magnetic Zeeman fields and Rashba-spin-orbit-coupling (RSO). The resulting approximately $10000$ magnetic configurations are too diverse to be reliably classified by a simple algorithm. We therefore use contrastive learning [33, 34], a recent unsupervised machine learning technique, which can be adapted to the physical symmetries. While contrastive learning was used to classify phases in the Ising, Compass, and Su-Schrieffer-Heeger models [35] using the schema of Ref. [33], we used Contrastive Clustering [36] to obtain the phases directly from the network. Our results show that this technique is highly effective for the classification of complex phase spaces of magnetic systems on SC.
@@ II.3 Heisenberg Parameters
The off-diagonal elements can be mapped to DMI-vectors, which are mostly lying in the xy-plane, i.e. the $D_{xy}$ component is very small, and are oriented perpendicular to the vector between the respective sites.
We also find that $B_{z}$ in the tight-binding model directly translates to a ${\boldsymbol{B}}$-field in $z$-direction in the classical model, as expected. In Fig. 3, the relative strength of the $4$-spin interaction is shown, which is calculated as the sum of the absolute values of all $C$ divided by the sum of the absolute values of all fitted parameters in Eq.(2). We found a small, but still significant contribution of $4$-spin interaction in well-fitted regions. This indicates that this simple seeming tight-binding system hosts complex multi-spin interactions beyond 2-spin interactions. The largest $4$-spin contribution is found in a region, where the fitting quality is not sufficient. Because the addition of $4$-spin interactions improves the fitting quality in that region, we speculate that this region is dominated by higher order or further reaching multi-spin interactions.
@@ II.4 Classification of Magnetic Phases using Contrastive Clustering
To better understand the ground state of the Heisenberg model with $B$-field and $4$-spin interactions described in Eq. (2), we employ Contrastive Clustering [36] an unsupervised learning technique where the learning of a latent representation and cluster assignment are performed simultaneously by comparing different samples. The samples get encoded into points in a representation space, the latent space. The goal is to attract points in this space if they correspond to the same phase and repel them if they come from different phases. To achieve this, we apply symmetry transformations to magnetic configurations. Two transformed images attract each other if they originate from the same configuration, else they repel, as visualized in Fig. 4. The details of the implementation of Contrastive Clustering follow Ref. [36]. We use Resnet18 [44] as the encoder network, Adam [45] to train the network, and train our network for 2000 cycles over the whole dataset ensuring sufficient convergence of our loss function. To deal with the unbalanced nature of our dataset, we adapted an oscillating computational temperature in the contrastive loss, as this has been shown to perform better with long-tail data [46], i.e., data where a small part of classes have a large number of sample points but the others are associated with only a few samples [47]. We implement the best-performing parameters of Ref. [46]. For the transformations, we use intrinsic symmetries of the magnetic structure identification, which are a rotation of the whole lattice by either 0°, 90°, 180°, or 270°, a shift of the lattice by a random amount in the x and y directions, and rotation of all spins together in a random direction. This last step violates the axial anisotropy of the system when the Rashba-spin-orbit coupling $\alpha$ or the magnetic field ${\boldsymbol{B}}$ is non-zero. We successfully classify the unlabeled spin data into different phases. Hence, we create a phase diagram without having to specify the phases in advance, which allows the detection of unexpected phases.
@@
We categorize the magnetic phases as follows. First, the artificial neural network splits the magnetic ground states into 40 clusters. Then, we reduce the number of clusters manually to $11$ magnetic phases, and one residual phase that contains all clusters with mixed magnetic configurations. Three initially mixed clusters were separated by adding an additional condition. For one of those, all states with $\mu<-2.0t$ belong to a magnetic 2x2-pattern, while the rest belong to AFM spin spirals. The other two contain harmonic spin spirals and 2q-spin-spirals but for $\alpha=0.0$ only contain harmonic spin spirals. Additionally, the artificial neural network (ANN) often combines skyrmion lattices and spin spirals with a periodicity that is similar to the skyrmion size in one cluster. To identify skyrmionic phases, we calculate the skyrmion number $N_{sk}$, see Eq. (4) and assign everything with $|N_{Sk}|>0.95$ as part of a skyrmionic phase, giving some tolerance for numerical precision.
@@ III.1 Vanishing Rashba-Spin-Orbit Coupling
We calculate the magnetic ground states of a $32\times 32$ system with periodic boundary conditions in real space. The ANN-classified magnetic state as described in Sec. II.4, in dependence on the magnetic coupling $J^{tb}$ and the chemical potential $\mu$ are shown in Fig. 6 for $\Delta=0.0$ and $\Delta=0.5t$ without RSO, i.e., $\alpha=0$. Fig. 6 a and b show phase diagrams for $\Delta=0.0$ and $\Delta=0.5t$, respectively. Examples of ground states are presented in Fig. 6 c.
@@
where ${\boldsymbol{M}}$ is the local magnetization and where we calculate the derivatives of the continuum extrapolation by the inverse Discrete Fourier Transform (DFT) of the product of the DFT frequencies and the DFT of the spins following $f^{\prime}(x)=\frac{1}{N}\sum_{k}e^{\mathrm{i}kx}(\mathrm{i}k)\sum_{z}e^{-\mathrm{i}kz}f(z)$. This avoids numerical artifacts for anti-ferromagnetic states that can appear with lattice adapted invariants [50] for the cost of deviations from the integer quantization of the skyrmion number. In most skyrmionic cases, we observe a skyrmion lattice, while in some cases singular skyrmions appear in a ferromagnetic background. In 2q-spin-spirals, we observe parallel spins in one direction and following a pattern of two alternating relative angles $\theta_{1}$ and $\theta_{2}$ in the perpendicular direction. For $\alpha=0$, these two angles average to $(\theta_{1}+\theta_{2})/2=\pi/2\pm 0.01$ in $85\%$ of cases. While in the harmonic FM and AFM spin spirals magnetization rotates with a wave-vector q along high symmetry crystallographic directions, the multi-q-spin-structures are superpositions of several rotations with high symmetry q-vectors [1]. Spin spirals and 2q-spin-spirals denote that the ANN cannot differentiate between 2q-spin-spirals and harmonic spin spirals, which only occurs for $\alpha=0.2$. Mixed denotes that the ANN created a label which we could not assign a definite magnetic state to.
Before discussing the influence of superconductivity, we start with $\Delta=0$, i.e., vanishing superconductivity. For large $J^{tb}>5.0t$, we only observe ferromagnetic (FM) and anti-ferromagnetic (AFM) phases. For smaller $J^{tb}$, we observe rich magnetic phases between the FM (A) and AFM (D) phase shown in Fig. 6b. For chemical potentials $-1.7t<\mu<-0.9t$, we find AFM spirals (F). At $\mu<-1.7t$, this phase transitions into row-wise AFMs (E). For lower chemical potentials, we find harmonic spin spirals with varying propagation directions other than the x- or y-axis (C). Lowering the chemical potential further results in x-y-row-wise spin spirals (B) and then in an FM phase. For $J^{tb}>1.2t$, the row-wise AFM phase first transitions into a 2x2 pattern (H) and then into the FM phase. Additionally around $\mu=-3.0t$ and $J^{tb}<0.4t$, there is a small region in parameter space with 2q-spin-spirals (J).
@@ III.2 Non-Vanishing Rashba-Spin-Orbit Coupling
Figure 8: Magnetic ground states in systems with Rashba-Spin-Orbit coupling and an additional out-of-plane magnetic field [$\alpha=0.2t$ and $B_{z}=0.02t$ ] and (a) $\Delta=0.0$, (b) $\Delta=0.5t$ with respect to $\mu$ and $J^{tb}$. The color white denotes regions with largely unoccupied bands. The hatched area leaves the validity regime of the model (see Fig. 3).
@@
The effect of RSO and a magnetic field on the other phases is displayed in Fig. 8. Besides the appearance of a new skyrmionic and spin spiral phase for low $\mu$ within the FM phase, the overall shape of the phases remains largely unchanged, but the magnetic structures within the phases change. The previous checkerboard AFM phase becomes a checkerboard AFM in superposition with a spin spiral of a characteristic angle of $0.124\pi$, which we called Rashba AFM here. Most parts of the FM phase remain ferromagnetic, but towards lower $\mu$, an additional skyrmionic and spin spiral phase occurs. The inner parts of the row-wise AFM phase is replaced by spin spirals, of which only some are AFM spirals. The 2x2 pattern does not occur anymore. For $\Delta=0.5t$, the 2q-spin-spiral phase still exists, but the ANN fails to differentiate it from harmonic row-wise spin spirals.
@@ IV Conclusions
Our results show that a simple effective model for magnetic moments coupled to a superconducting substrate with Rashba spin-orbit coupling accommodates an unexpectedly rich phase diagram, helping with the search for exotic and noncollinear phases at magnetic surfaces, which can be effectively categorizes by an ANN with contrastive learning algorithm. The combination of the 2D tight-binding model with Monte Carlo simulations permits us to comprehensively categorize the magnetic ground states in the SC and non-SC regimes. In addition to the collinear FM and AFM configurations as well as harmonic noncollinear states known from one-dimensional models, 2D systems accommodate manifold multi-q noncollinear structures and non-trivial collinear configurations like Rashba-AFM or 2q-spin-spirals. An interesting property of an ultrathin film on a square lattice is the appearance of multi-spin interactions that have been predicted to exist in itinerant magnets but are only important in few experimental studies so-far like the nanoskyrmion lattice in Fe/Ir(111) [2]. Our phase diagrams help researchers working with ultrathin magnetic films, showing what types of magnetic ordering, they can expect in their system or, vice versa, to estimate model parameters from the observed magnetic ordering. Generally, superconductivity decreases the area of the FM phase in favor of AFM and non-collinear structures, while the Rashba coupling favors noncollinearity.

### v2

@@ I Introduction
There is a vast versatility of magnetic phases in the physics of interfaces. Among ferromagnetic and antiferromagnetic spin orientations in the ground state, recently noncollinear phases including spirals, multi-q states [1, 2, 3], canted AFMs [4], bubbles, and topological magnetism including skyrmions [5] and merons [6], as well as further exotic phases like altermagnets [7] have been theoretically and experimentally discovered. Magnetic phases form a cornerstone of today’s and future storage technology [8, 9, 10] and are fundamentally interesting because of exotic classical and quantum quasiparticle excitations that could be hosted in frustrated or correlated magnets [11, 12]. For modeling purposes and for detecting yet novel magnetic phases, the question arises which conceptionally simple models host a large number of magnetic phases that can be tuned by experimentally accessible parameters. Particularly interesting in this regard are systems combining magnetism and superconductivity. While numerous studies of one-dimensional magnetic systems in proximity to s-wave superconductors exist, motivated by the interest in Majorana physics [13, 14, 15, 16], the data on phase diagrams of their two-dimensional counterparts have been very limited up to now [17, 18]. Ground state studies on quantum and classical magnetic moments coupled to itinerant electrons found evidence of chiral order [19], topological electronic order [20, 21], and multi-spin interactions [22] in frustrated triangular lattices without superconductivity. Motivated by recent studies on superconducting 1D systems [13], we therefore consider a two-dimensional square lattice tight-binding model with local magnetic moments, Rashba spin-orbit coupling, and superconductivity. Our model is effectively describing localized classical magnetic moments that are coupled by the itinerant electrons of a surface state, not specifically bound to any parameter regime. The knowledge of possible ground states in those complex structures is required for several reasons. First, exotic topological states have been found in 2D itinerant magnets in proximity to a superconductor (SC) [17, 18]. Second, interesting multi-spin interactions can be present in such systems [23, 1] that might lead to unexpected magnetic and electronic phases. Due to big datasets in the study of magnetic systems, machine learning has been applied extensively in this field, which we adopt as well. Corresponding techniques have especially been employed to predict the parameters of the magnetic Hamiltonian [24], to retrieve the topological charge from dynamic chiral magnetic structures [25], and to construct phase diagrams for skyrmionic systems [26, 27]. Additionally, unsupervised learning has found a variety of applications in the field of microscopic magnetism [28, 29, 30] and the identification of topological phases [31, 32].
@@
In this manuscript, we analyze the ground states of magnetic moments coupled by a spinful tight-binding model with proximity induced s-wave superconductivity, local magnetic Zeeman moments and Rashba-spin-orbit-coupling (RSO). The resulting approximately $10000$ magnetic configurations are too diverse to be reliably classified by a simple algorithm. We therefore use contrastive learning [33, 34], a recent unsupervised machine learning technique, which can be adapted to the physical symmetries. While contrastive learning was used to classify phases in the Ising, Compass, and Su-Schrieffer-Heeger models [35] using the schema of Ref. [33], we used Contrastive Clustering [36] to obtain the phases directly from the network. Our results show that this technique is highly effective for the classification of complex phase spaces of magnetic systems on SC.
@@ II.3 Heisenberg Parameters
The off-diagonal elements can be mapped to DMI-vectors, which are mostly lying in the xy-plane, i.e. the $D_{xy}$ component is very small, and are oriented perpendicular to the vector between the respective sites. They obey the system’s rotational and translational lattice symmetries.
We also find that $B_{z}$ in the tight-binding model directly translates to a ${\boldsymbol{B}}$-field in $z$-direction in the classical model, as expected. In Fig. 3, the relative strength of the $4$-spin interaction is shown, which is calculated as the sum of the absolute values of all $C$ divided by the sum of the absolute values of all fitted parameters in Eq.(2). We found a small, but still significant contribution of $4$-spin interaction in well-fitted regions. This indicates that this simple seeming tight-binding system hosts complex multi-spin interactions beyond $2$-spin interactions. The largest $4$-spin contribution is found in a region, where the fitting quality is not sufficient. Because the addition of $4$-spin interactions improves the fitting quality in that region, we speculate that this region is dominated by higher order or further reaching multi-spin interactions. Adding non-vanishing RSO of $\alpha=0.2t$ causes only minimal changes in the pattern of $4$-spin interactions. The relative strength is lowered by $\sim 28\%$ on average, mostly due to an absolute increase in the strength of $2$-spin interactions due to the addition of DMI.
@@ II.4 Classification of Magnetic Phases using Contrastive Clustering
To better understand the ground state of the Heisenberg model with $B$-field and $4$-spin interactions described in Eq. (2), we employ Contrastive Clustering [36] an unsupervised learning technique where the learning of a latent representation and cluster assignment are performed simultaneously by comparing different samples. The samples get encoded into points in a representation space, the latent space. The goal is to attract points in this space if they correspond to the same phase and repel them if they come from different phases. To achieve this, we apply symmetry transformations to magnetic configurations. Two transformed images attract each other if they originate from the same configuration, else they repel, as visualized in Fig. 4. The details of the implementation of Contrastive Clustering follow Ref. [36]. We use Resnet18 [44] as the encoder network, Adam [45] to train the network, and train our network for 2000 cycles over the whole dataset ensuring sufficient convergence of our loss function. The output of Resnet18 is a 512-dimensional vector. On top of that, two multilayer perceptrons are applied as projection heads, which project the output of the encoder to a 128-dimensional instance space (visualized in Fig. 5) and a 40-dimensional clustering space. They both consist of one hidden layer with 512 neurons and a rectified linear unit (ReLU) as activation function. To deal with the unbalanced nature of our dataset, we adapted an oscillating computational temperature in the contrastive loss, as this has been shown to perform better with long-tail data [46], i.e., data where a small part of classes have a large number of sample points but the others are associated with only a few samples [47]. We implement the best-performing parameters of Ref. [46]. For the transformations, we use intrinsic symmetries of the magnetic structure identification, which are a rotation of the whole lattice by either 0°, 90°, 180°, or 270°, a shift of the lattice by a random amount in the x and y directions, and rotation of all spins together in a random direction. This last step violates the axial anisotropy of the system when the Rashba-spin-orbit coupling $\alpha$ or the magnetic field ${\boldsymbol{B}}$ is non-zero. We successfully classify the unlabeled spin data into different phases. Hence, we create a phase diagram without having to specify the phases in advance, which allows the detection of unexpected phases.
@@
We categorize the magnetic phases as follows. First, the artificial neural network splits the magnetic ground states into $40$ clusters. Often, multiple clusters belong to the same magnetic phase, e.g., there are multiple clusters with harmonic spin spirals but different ranges of relative angles between the nearest magnetic moments. Those clusters are then manually combined to $11$ magnetic phases, which we identify by spot checks. The number of clusters is chosen larger than the number of expected physical phases to avoid misclassification. For example, when limited by too few clusters, the ANN could assign an $87^{\circ}$-spiral to an antiferromagnetic cluster instead of a spiral one. When a cluster cannot be clearly assigned to a single magnetic phase, i.e., we identify different magnetic phases within one cluster, we label that cluster as mixed. Three clusters initially labeled as mixed could be separated by adding an additional condition. For one of those, all states with $\mu<-2.0t$ belong to a magnetic 2x2-pattern, while the rest belong to AFM spin spirals. The other two contain harmonic spin spirals and 2q-spin-spirals but for $\alpha=0.0$ only contain harmonic spin spirals. Additionally, the artificial neural network (ANN) often combines skyrmion lattices and spin spirals with a periodicity that is similar to the skyrmion size into one cluster. To identify skyrmionic phases, we calculate the skyrmion number $N_{sk}$, see Eq. (4) and assign everything with $|N_{Sk}|>0.95$ as part of a skyrmionic phase, giving some tolerance for numerical precision.
@@ III.1 Vanishing Rashba-Spin-Orbit Coupling
We calculate the magnetic ground states of a $32\times 32$ system with periodic boundary conditions in real space. The ANN-classified magnetic state as described in Sec. II.4, in dependence on the magnetic coupling $J^{tb}$ and the chemical potential $\mu$ are shown in Fig. 6 for $\Delta=0.0$ and $\Delta=0.5t$ without RSO, i.e., $\alpha=0$. Fig. 6 b and c show phase diagrams for $\Delta=0.0$ and $\Delta=0.5t$, respectively. Examples of ground states are presented in Fig. 6 a.
@@
where ${\boldsymbol{M}}$ is the local magnetization and where we calculate the derivatives of the continuum extrapolation by the inverse Discrete Fourier Transform (DFT) of the product of the DFT frequencies and the DFT of the spins following $f^{\prime}(x)=\frac{1}{N}\sum_{k}e^{\mathrm{i}kx}(\mathrm{i}k)\sum_{z}e^{-\mathrm{i}kz}f(z)$. This avoids numerical artifacts for antiferromagnetic states that can appear with lattice adapted invariants [50] for the cost of deviations from the integer quantization of the skyrmion number. In most skyrmionic cases, we observe a skyrmion lattice, while in some cases singular skyrmions appear in a ferromagnetic background. In 2q-spin-spirals, we observe parallel spins in one direction and following a pattern of two alternating relative angles $\theta_{1}$ and $\theta_{2}$ in the perpendicular direction. For $\alpha=0$, these two angles average to $(\theta_{1}+\theta_{2})/2=\pi/2\pm 0.01$ in $85\%$ of cases. While in the harmonic FM and AFM spin spirals magnetization rotates with a wave-vector q along high symmetry crystallographic directions, the multi-q-spin-structures are superpositions of several rotations with high symmetry q-vectors [1]. Spin spirals and 2q-spin-spirals denote that the ANN cannot differentiate between 2q-spin-spirals and harmonic spin spirals, which only occurs for $\alpha=0.2$. Mixed denotes that the ANN created a label which we could not assign a definite magnetic state to.
Before discussing the influence of superconductivity, we start with $\Delta=0$, i.e., vanishing superconductivity. For large $J^{tb}>5.0t$, we only observe ferromagnetic (FM) and antiferromagnetic (AFM) phases. For smaller $J^{tb}$, we observe rich magnetic phases between the FM (A) and AFM (D) phase shown in Fig. 6b. For chemical potentials $-1.7t<\mu<-0.9t$, we find AFM spirals (F). At $\mu<-1.7t$, this phase transitions into row-wise AFMs (E). For lower chemical potentials, we find harmonic spin spirals with varying propagation directions other than the x- or y-axis (C). Lowering the chemical potential further results in x-y-row-wise spin spirals (B) and then in an FM phase. For $J^{tb}>1.2t$, the row-wise AFM phase first transitions into a 2x2 pattern (H) and then into the FM phase. Additionally around $\mu=-3.0t$ and $J^{tb}<0.4t$, there is a small region in parameter space with 2q-spin-spirals (J).
@@ III.2 Non-Vanishing Rashba-Spin-Orbit Coupling
Figure 8: Magnetic ground states in systems with Rashba spin-orbit coupling $\alpha=0.2t$, an additional out-of-plane magnetic field $B_{z}=0.02t$ and (a) $\Delta=0.0$, (b) $\Delta=0.5t$ with respect to $\mu$ and $J^{tb}$. The color white denotes regions with largely unoccupied bands. The hatched area leaves the validity regime of the model (see Fig. 3).
@@
The effect of RSO and a magnetic field on the other phases is displayed in Fig. 8. Besides the appearance of a new skyrmionic and spin spiral phase for low $\mu$ within the FM phase, the overall shape of the phases remains largely unchanged, but the magnetic structures within the phases change. The previous checkerboard AFM phase becomes a checkerboard AFM in superposition with a spin spiral of a characteristic angle of $0.124\pi$, which we called Rashba AFM here. Most parts of the FM phase remain ferromagnetic, but towards lower $\mu$, an additional skyrmionic and spin spiral phase occurs. The inner parts of the row-wise AFM phase is replaced by spin spirals, of which only some are AFM spirals. The 2x2 pattern does not occur anymore. For $\Delta=0.5t$, the 2q-spin-spiral phase still exists, but the ANN fails to differentiate it from harmonic row-wise spin spirals. We also observe a larger area in the phase diagram labelled B/C/E/F, where multiple phases coexist. When inspecting single samples in this area, we find that the phases are accurately assigned by the ANN and that this apparent disorder truly reflects the large diversity of magnetic configurations in that area.
@@ IV Conclusions
Our results show that a simple effective model for magnetic moments coupled to a superconducting substrate with Rashba spin-orbit coupling accommodates an unexpectedly rich phase diagram, helping with the search for exotic and noncollinear phases at magnetic surfaces, which can be effectively categorized by an ANN with contrastive learning algorithm. The combination of the 2D tight-binding model with Monte Carlo simulations permits us to comprehensively categorize the magnetic ground states in the SC and non-SC regimes. In addition to the collinear FM and AFM configurations as well as harmonic noncollinear states known from one-dimensional models, 2D systems accommodate manifold multi-q noncollinear structures and non-trivial collinear configurations like Rashba-AFM or 2q-spin-spirals. An interesting property of an ultrathin film on a square lattice is the appearance of multi-spin interactions that have been predicted to exist in itinerant magnets but are only important in few experimental studies so-far like the nanoskyrmion lattice in Fe/Ir(111) [2]. Our phase diagrams help researchers working with ultrathin magnetic films, showing what types of magnetic ordering, they can expect in their system or, vice versa, to estimate model parameters from the observed magnetic ordering. Generally, superconductivity decreases the area of the FM phase in favor of AFM and non-collinear structures, while the Rashba coupling favors noncollinearity.